OpenStudy (anonymous):

Ben dives from a 36-foot platform. The equation h=-16t^2+14t+36 models the dive. How long will it take Ben to reach the water?

OpenStudy (whpalmer4):

Let's look at that equation: \[h(t) = -16t^2+14t+36\]At \(t=0\) he hasn't yet taken his dive: \[h(0) = -16(0)^2 + 14(0) + 36 = 0 + 0 + 36\]36 feet up, and it's a 36 foot tower. What will the value of \(h(t)\) be when he reaches the water?

OpenStudy (anonymous):

Zero?

OpenStudy (anonymous):

Wait would it be 72?

OpenStudy (whpalmer4):

Yes, the water is at 0. If you sketch the graph of that function, it starts out at 36 at t=0 and gets smaller: t=0 is the peak. So if the water is at height 0, set \(h(t) = 0 = -16t^2 + 14t + 36\] and solve for the value of \(t\) that makes that true.

OpenStudy (anonymous):

put h=36 and find the values of t, \[36=-16t ^{2}+14t+36\] \[-2t(-8t+7)=0,t=0`gives~the~time~of~starting,-8t+7=0,t=\frac{ 7 }{ 8 }\]

OpenStudy (anonymous):

Oh wow thank you guys so much!!!!

OpenStudy (anonymous):

yw

OpenStudy (whpalmer4):

No, that's incorrect!

OpenStudy (whpalmer4):

The water is at height h = 0! Do as I said and solve \[h(t) = 0 = -16t^2+14t +36 \]for the (positive) value of \(t\) where \(h(t) = 0\). If you think the other approach is correct, then consider this: At t = 7/8, look at the height of the diver: \[h(\frac{7}{8}) = -16(\frac{7}8)^2+14(\frac{7}8)+36 = -\frac{49}{4} + 49/4+36 = 36 \]That's not water level, that's the level at which he started!

OpenStudy (whpalmer4):

This is why we check our answers :-)

OpenStudy (anonymous):

sorry i am wrong Mr .whpalmer is correct.  i took h from top, actually it is from the ground. when t=0, h=36 thank you for correcting me.

OpenStudy (whpalmer4):

 Here's the path of the diver expressed as height vs. time. Notice that he goes up initially and then down. What @surjithayer found was the time at which the diver returned to his initial height. An equally reasonable thing for a problem to ask, but not what this one did. Whenever I do a word problem like this, I like to reread the problem before I check my answer, just to be certain that I've actually solved the same problem!